Kewin has $2.15 in dimes and quarter. He has 3 less dime than quarters. How many of each does he have?
Answer by checkley75(3666)

25q+10(q-3)=215
25q+10q-30=215
35q=215+30
35q=245
q=245/35
q=7 number of quarters.
7-3=4 number of dimes.
proof
25*7+10*4=215
175+40=215
215=215